TITLE: Calculating the probability of a coin falling on its side
QUESTION [10 upvotes]: A classical example that's given for probability exercises is coin flipping. Generally it is accepted that there are two possible outcomes which are heads or tails. However, it is possible in the real world for a coin to also fall on its side which makes a third event ( $P(\text{side}) = 1 - P(\text{heads}) - P(\text{tails})$ ?). How would a mathematician go about calculating the probability of that? Could it be done by simply using respective surface areas or would a proper model be more complex?

REPLY [2 votes]: While I am and always have been a big proponent of the edge factoring into probability in just a quirky enough way as to disrupt an absolute $50-50$ "fair" game and create the equivalent to the least profitable house advantage in existence kind of like $0$ and $00$ on a roulette wheel, the idea that the probability is as good as $1$ in $6000$ or $.0166%$ seems kind of high even for a nickel.
I guess I think about it in terms of if possibility exists in any way, regardless of probability, in the moment of the flip is the coin in a state of fluctuation that is suggestive of one of $2$ outcomes but indeed it is not a definite zero-sum game in that a $3$rd party, as moronic as his speculation may be, could be the benefactor of the outcome should he place his position on a "house" win... I dunno, it's that kind of stuff that distracts me from learning anything productive! :)